OpenStudy (anonymous):

what is 1/3 times 200 plus 1/10 please explain the steps

OpenStudy (anonymous):

1/3 * 200 = 200/3 +1/10 you must find a common denominator which is 30 in this case so you multiply the top and bottom until you get 30 in the denominator and it becomes 2003/30

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

is their an answer

OpenStudy (anonymous):

Here is how\[\frac{ 1 }{ 3 }\times200+\frac{ 1 }{ 10 }\]\[\frac{ 2000 }{ 30 }+\frac{ 3 }{ 30 }\]\[\frac{ 2003 }{ 30 }\]

OpenStudy (anonymous):

200/3 + 1/10 common denominator is 30 so do, 200 * 10 and 3* 10, and 1*3 and 10 * 3

OpenStudy (anonymous):

yes i posted it 2003/30

OpenStudy (anonymous):

2000/30+3/30 = 2003/30

OpenStudy (anonymous):

there you go ^

OpenStudy (anonymous):

@Zekarias you did not show enough steps

OpenStudy (anonymous):

the answer is 66 and 23/30

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

it is the same as 2003/30

OpenStudy (anonymous):

I know i just realized that sorry

OpenStudy (anonymous):